OpenStudy (anonymous):

find the magnitude of wx for w(-3,5,4) and x(9,5,3) i think its sqrt145

OpenStudy (anonymous):

That looks right, but I'm not for sure. Sorry

OpenStudy (anonymous):

thanks anyways

OpenStudy (some.random.cool.kid):

multiply

jimthompson5910 (jim_thompson5910):

wx = (9-(-3), 5-5, 3-4) wx = (12, 0, -1) |wx| = sqrt(12^2 + 0^2 + (-1)^2) |wx| = sqrt(145) so you have the right answer

OpenStudy (anonymous):

thanks